A 45kg block ar rest was pushed by a 280N horizontal force up a 30degree inclined plane with a roughness coefficient of 0.15, what is the netforce acting on the block


1
Expert's answer
2021-12-02T10:04:37-0500


1) Determine the normal force:


"N=mg\cos\theta+F\sin\theta."

2) Find the force of friction:


"f=\mu N=\mu(mg\cos\theta+F\sin\theta)."

3) The net force:


"F_\text{net}=F\cos\theta-f,\\\space\\
F_\text{net}=F\cos\theta-\mu(mg\cos\theta+F\sin\theta)=164\text{ N}."